hi everyone. I'm trying to find the formula for this question, i've looked amongst my notes and I just don't have it. I'm trying to find the principle brian requires 75,000 5years from now, he is offered an interest rate of 5% pa compounded annually. how much would he need to invest now in order to satisfy his requirement?
do I have the right formula? P=I/RT? 75,000/0.05x5 = 30,000?
a steer in the right direction would be so amazing. thank you

Answer by Theo(13342)

the formula is f = p * (1+r)^n

f is the future value
p is the present value
r is the interest rate per time period
n is the number of time periods.

in your problem:

f = 75,000
p = what you want to find.
r = .05 per year
n= 5 years

your formula becomes:

75,000 = p * 1.05^5

divide both sides of this equation by 1.05^5 to get:

75,000 / 1.05^5 = p

solve for p to get p = 58764.46249

i believe you are using the simple interest formula and not the compound interest formula that i showed you above.

the simple interest formula is i = prt

i is the interest
p is the principal, otherwise known as the present value.
r is the interest rate per time period
t is the number of time periods.

not the right formula to use for compound interest type problems.
additionally, you were given the future value, not the present value.

to find the future value with a simple interest type formula, you would use the following formula:

f = p + i = p + prt

doesn't matter.
still the wrong formula to use since your problem is a compound interest rate type problem and not a simple interest rate type problem.